### v1

# Readout electronics for low occupancy High-Pressure Gas TPCs

Abstract
High-Pressure Gas Time Projection Chambers have benefits such as low energy threshold, magnetisability, and acceptance, making them ideal for neutrino experiments such as DUNE. We present the design of an FPGA-based solution optimised for Gaseous Argon Near Detector (ND-GAr), which is part of the Phase-II more capable near detector for DUNE. These electronics reduce the cost significantly compared to using collider readout electronics which are typically designed for much higher occupancy and therefore, for example, need much larger numbers of FPGAs and power per channel. We demonstrate the performance of our electronics with the Teststand for an Overpressurised Argon Detector (TOAD) at Fermilab in the US at a range of pressures and gas mixtures up to , reading out channels from a multi-wire proportional chamber. The operation took place between April and July of 2024. We measure the noise characteristics of the system to be sufficiently low and we identify sources of noise that can be further mitigated in the next iteration. We also note that the cooling scheme used in the test requires improvement before full-scale deployment. Despite these necessary improvements, we show that the system can fulfil the needs of a HPgTPC for a fraction of the price of collider readout electronics.
- TPC
- Time Projection Chamber
- FEC
- Front-End Card
- DFT
- Discrete Fourier Transformation
- OROC
- Outer Read-Out Chamber
- ADC
- Analogue to Digital Converter
- RMS
- Root Mean Square
- TOAD
- Teststand for an Overpressurised Argon Detector
- DUNE
- Deep Underground Neutrino Experiment
- HPgTPC
- High-Pressure Gas Time Projection Chamber
- ND-GAr
- Gaseous Argon Near Detector
- FPGA
- Field Programmable Gate Array
- DC/DC
- Direct Current to Direct Current
- IDC
- Insulation-Displacement Contact
- PAT
- Power, Aggregation and Timing
- ROC
- Read-Out Chamber
- DAQ
- Data Acquisition
- LDO
- Low-Dropout regulator
- PoE
- Power over Ethernet
- DC
- Direct Current
- HV
- High Voltage
- IROC
- Inner Read-Out Chamber
- SFP
- Small Form-Factor Pluggable
- CCM
- Control Configuration and Monitoring
- PCB
- Printed Circuit Board
- LVDS
- Low-Voltage Differential Signalling
- FSM
- Finite State Machine
- I2C
- Inter-Integrated Circuit
- SPI
- Serial Peripheral Interface
- PMbus
- Power-Management Bus
- UDP
- User Datagram Protocol
- TCP
- Transmission Control Protocol
- RR
- Round-Robin
- LHC
- Large Hadron Collider
- LAr
- Liquid Argon
- K-S
- Kolmogorov-Smirnov
- CLI
- Command Line Interface
- ASIC
- Application Specific Integrated Circuit
- MWPC
- Multi-Wire Proportional Chamber
- GEM
- Gas Electron Multiplier
- ECal
- Electromagnetic Calorimeter
- COTS
- Commercial Off the Shelf
- ND
- near detector
- ALICE
- A Large Ion Collider Experiment
1 Introduction
Gas detectors offer greatly reduced energy thresholds, magnetisability, and 4 acceptance, which are not currently achieved with other detector technologies [1]. However, gases have much lower density than solid or liquid detectors, reducing the interaction rate for particles passing through. The advent of MW-power neutrino beams means that detectors with gas targets will, for the first time, be practical for long-baseline neutrino experiments. Despite these higher beam powers, a gas Time Projection Chamber (TPC) detector in these beams would still need to be pressurised to accumulate a sufficient number of interactions.
Harnessing the capabilities of such a High-Pressure Gas Time Projection Chamber (HPgTPC) for the Deep Underground Neutrino Experiment (DUNE) would provide important insights into neutrino-argon interactions and enable key measurements of the acceptance corrections of other detectors. As such, a gaseous argon TPC surrounded by an Electromagnetic Calorimeter (ECal) has been recognised as crucial to the experiment’s physics program by the recent US P5 process [2], and the DUNE collaboration’s Phase II white paper [3].
The detector’s position resolution is a critical factor in achieving the required momentum resolution and detection thresholds for complex low-energy interactions with high multiplicity. The resolution obtainable is closely related to the number of readout channels that the detector can accommodate and the sampling frequency of those channels. Whilst the numbers are not final, instrumenting a diameter TPC sufficiently to achieve the desired low energy thresholds for the DUNE would require channels with sampling frequency [1]. The development of readout electronics in the particle physics community has previously focused on the design and operation of high-throughput radiation-hard electronics required at the Large Hadron Collider (LHC)-based experiments [4]. Naively scaling such designs for neutrino experiments would be prohibitively expensive. For example, in the context of the DUNE Phase-II upgrade, adopting electronics like those used for sPHENIX [5] or the TPC of A Large Ion Collider Experiment (ALICE) [6] with its readout electronics requires a high-specification Field Programmable Gate Array (FPGA) to readout groups of 10 front-end cards [7], providing a resolution of 1- in the TPC drift direction. However, given the hundreds of thousands of channels required to meet DUNE’s resolution needs, this approach would result in a total cost larger than possible, given budgetary constraints for the DUNE near detector. In fact, the electronics were previously listed as the most expensive component for the Gaseous Argon Near Detector (ND-GAr) for DUNE.
Since neutrino experiments have, by their nature, much lower occupancy and hence lower throughput than LHC experiments, a re-optimisation of the readout system enables significant cost savings whilst retaining physics performance.
In the following sections we present the design of a scalable highly re-configurable FPGA based solution optimised to be affordable for the ND-GAr DUNE Phase-II detector [3]. This system delivers significant cost-saving compared to the systems described above while showing no performance degradation in an HPgTPC environment. A prototype has been produced and was then sent for testing to Fermilab as part of the Teststand for an Overpressurised Argon Detector (TOAD). DUNE has not decided on an amplification stage for its HPgTPC but a Gas Electron Multiplier (GEM)-based solution is the current front runner. Conversely, TOAD uses a Multi-Wire Proportional Chamber (MWPC) from the ALICE, due to its availability at the time of the test. This chamber, called an Outer Read-Out Chamber (OROC) [8], is deployed in the high-pressure vessel from the UK high-pressure facility. Large parts of this setup have previously been used to test gas mixtures for high-pressure TPCs [9].
2 Readout setup
A key challenge for future neutrino experiments is the harsh conditions in which the electronics must operate. In contrast to collider experiments, those conditions are not due to extensive radiation but instead factors such as cryogenic temperatures present in Liquid Argon (LAr)-based detectors and high pressure such as in an HPgTPC based detector. In the case of the high-pressure detectors an important consideration is the number of feedthroughs required to power and communicate with the electronics, so there is a need to aggregate data inside the vessel. Whilst power dissipation and thermal management present additional challenges, they can be addressed using water cooling through a small number of vessel penetrations. Hence, a readout system will need to optimise the number of cables required to stream data from the large amount of analogue readout channels.
The proposed solution outlined in this paper uses a fast-sampling Front-End Card (FEC) with zero-suppression capability and a custom FPGA-based Power, Aggregation and Timing (PAT) card, both residing in-situ within the high-pressure volume. Data are then streamed over a standard Ethernet-based network to the DUNE Data Acquisition (DAQ) software operating on standard servers, avoiding extensive requirements within the pressurised environment. The block diagram of the proposed readout system is shown in figure 1.
Compared to a collider experiment, the channels in neutrino experiments have very low occupancy (in the similarly sized ALICE, collisions occur at and each generate thousands of tracks, whereas in DUNE, beam spills happen at and generate only tens of tracks per spill). This difference means far less processing is required as there are many fewer above threshold hits. Therefore, the system can have fewer FPGAs per channel (1 per channels with the particular FECs in this design) and use much lower specification FPGAs as we stream out all hits to a software-based DAQ. Furthermore, the lower data rates from the FEC allow the native transceivers on the SAMPAs to be used to communicate with the PAT so we also do not need the additional transceiver modules that are required for higher occupancy detectors [7]. These differences result in a much lower cost to instrument the hundreds of thousands of channels necessary for a neutrino HPgTPC. Even though it is a prototype system manufactured at small scale, the readout for TOAD (including FECs, PAT cards and all the hardware to connect and power them) cost less than $ for channels. Scaling this up to the full DUNE Phase-II near detector (ND), ignoring likely economies of scale, gives a conservative cost estimate of $1.4 million.
2.1 Hardware layout
As described, our system requires two types of custom hardware boards. The first type are the FEC boards, which host the digitiser Application Specific Integrated Circuits and connect to the detector itself. The second type are the PAT boards, which utilise an FPGA device for interfacing and aggregating the data for a multitude of FECs as well as providing them with power and distributing clock signals. Both the FECs and the PATs are custom designed, with the FECs designed for the TOAD test beam detector and its ALICE OROC. The flexibility of the FPGA-based design of the PAT cards means that the PAT can be reused even if the FECs change, so, whilst designed for TOAD, this board could be used with minimal modifications for the final ND-GAr detector or other detectors with similar data volume and occupancy. In designing the hardware, we avoided components with known air gaps e.g. lidded FPGAs packages. However, as there is no standard for electronics operation in the regime, a goal of TOAD is to test the operation of components at high pressures.
2.1.1 Front-End Card
GEMs are currently the leading candidate for the charge readout for ND-GAr. The SAMPA ASIC used for this readout system was developed by the ALICE collaboration for the GEM upgrade of their TPC [7]. This upgraded detector has similar characteristics to those needed for ND-GAr, so SAMPA is expected to meet the requirements of ND-GAr. The parameters of the SAMPA measured by ALICE can be seen in table 1 [10, 7]. The electronics system described here is therefore designed around deploying the SAMPA ASIC [10] used by ALICE, in a high-pressure environment, and at a much lower cost than in the full ALICE system.
| Specification | GEM |
| Supply Voltage | |
| Detector Capacitance | 40- |
| Peaking Time | |
| Shaping Order | 4th |
| Equivalent Noise | <e Cd= |
| Sampling Rate | , up to |
| Total Power (Idle) |
In addition, using an ASIC that has been proven to work at scale for another particle physics detector minimises risks and allows us to benefit from the extensive studies of their performance and limitations that have already been performed.
Another important reason for the choice of the SAMPA chips is their ability to sample at with fast Analogue to Digital Converters. Gas detectors have about an order of magnitude faster electron drift velocities than liquid; hence, this high sample frequency is necessary.
For TOAD, the ALICE MWPC has a group of three connectors per slot, which correspond to a total of 64 channels. Given that the SAMPA ASIC has only 32 channels, the FEC designed for TOAD, which can be seen in figure 2, has three connectors and hosts two SAMPAs to capture all 64 channels. The final ND-GAr amplification stages may have a different number of channels per connector, however the more SAMPAs per FEC the lower the cost of the system so designing for 2 SAMPAs represents a reasonable worst-case cost. To minimise the required readout channels connecting to the PAT board, the two SAMPAs on each FEC are daisy-chained as the low occupancy of neutrino experiments does not require the full bandwidth of the SAMPA readout link. However, as this is a prototype detector, the layout of the FEC has been designed to allow each SAMPA to be read out individually with minimal changes to the fitted components and no re-design of the Printed Circuit Board (PCB).
For TOAD, the FECs were operated in the configuration shown in table 2. This configuration gives an expected current draw of /channel with the FEC operating at full capacity. It is not expected that the FEC will have to operate at full power due to low occupancy and zero suppression. Whilst the design of the FEC is specific to the setup for TOAD, the operation at TOAD showcases the expected performance of SAMPA in a setup at an environment much closer to that required of a HPgTPC such as ND-GAr than previously tested.
| Parameter | Value |
| Supply Voltage | |
| Reference Clock | |
| Sampling Frequency | |
| Operation Mode | Streaming/Trigger |
2.1.2 Power Aggregation and Timing cards
The PAT card’s primary function is to interface between the FEC cards and a standard ethernet network where data can be streamed to servers for data selection and storage. The prototype, designed at Imperial College London, utilises the advantages of FPGAs as the main data processing unit for the PAT. An image of the prototype board can be seen in figure 3. The primary components on the PAT cards are listed in table 3.
| Component | Quantity |
| Kintex Ultrascale FPGA | |
| Insulation-Displacement Contact (IDC) connectors | |
| Clock Synthesizers | |
| Direct Current to Direct Current (DC/DC) Converters |
Data aggregation is a primary requirement that drives the design of the PAT. The cards are designed to take inputs from 56 FECs each streamed over a differential pair into the FPGA package. The received data payloads from the many FECs are then extracted and propagated through a multitude of buffers and Round-Robin (RR) logic to convert them into a single output stream. This output data stream is then wrapped into an Ethernet-based protocol and sent over the FPGA serial transceivers to the COTS network switch. The full layout can be seen in figure 4. Due to the low occupancy of the detector, the aggregation factor of 56 to 1 can be achieved easily within the FPGA fabric.
As well as data aggregation, the powering and timing roles of the PAT are also crucial. For powering, we investigated two possible solutions, for which a summary of key metrics is provided in table 4. The first solution provides to each PAT via dedicated terminals on the board. The FPGA DC/DC converter then outputs which is further fed into the FEC DC/DC converter to be transformed to the needed by the FEC on-board Low-Dropout regulator (LDO). The advantage of this scheme is that each PAT card receives a dedicated power line, minimising potential interference with signal traces on the board. The disadvantage is that ND-GAr will need dedicated power feedthroughs in the pressure vessel, which increases the overall feedthrough count.
The second ‘power over serial’ powering scheme is inspired by Power over Ethernet (PoE), and aims to allow PAT cards placed within the pressure vessel to receive power over the incoming network cable. The circuit diagram used can be seen in figure 5. However, some modifications to the normal PoE standard are required due to the intended environment. PoE works by injecting a Direct Current (DC) component on top of the alternating data component using transformers. Copying the same methodology for a magnetised detector wouldn’t be viable, as transformer performance may be impacted by the surrounding magnetic field. Therefore, an alternative system was developed using air-cored inductors to extract the differential signal from the bias voltage.
Such a scheme provides the following two advantages. Primarily, the number of feedthroughs and cables required for the complete detector readout would be reduced compared to the direct power solution. A second major difference with this option is that a board that is outside of the pressure vessel is required to receive these power plus data signals. This second board would allow additional buffering and processing to be done on the FPGA if necessary. In the explored design, both the PAT boards inside and outside of the pressure vessel were based on the same board design with minor differences in the fitted components.
The implementation on an early prototype for the PAT cards was found not to give sufficient signal stability, so further optimisations are required to achieve a production level system with this powering option. This work is underway with a new dedicated hardware board allowing modifications to both the power injection and the power decoupling circuits for further analysis. However, as it has not yet converged we therefore chose the first solution (direct power) for the board design discussed here.
In addition, a study, summarised in table 4, was done to estimate the number of feedthroughs needed for both options and also a third potential solution which would use optical fibres to stream out the data and a separate power line for each PAT. This option allows several PAT cards to share one output feedthrough as optical fibres are smaller in form factor than the electrical connectors needed.
| Component | Power over Serial | Direct Power | |
| Readout Type | Electrical | Optical | Electrical |
| # PAT (In-Pressure) | 196 | 196 | 196 |
| # PAT (Out-Pressure) | 22 | 0 | 0 |
| # FPGA | 218 | 196 | 196 |
| # Feedthrough | 196 | 76 | 206 |
| # Net. Switch Eth. Ports | 22 | 196 | 196 |
Decisions on the power distribution of the system also have an impact on the timing. For the Power over Serial method, timing can be provided to the upstream PAT card that is outside the pressure vessel and all the cards within the vessel can then be synchronised by the serial transceiver recovered clock. In the case of Direct Power, the data stream is separate to power and is running through commercial Small Form-Factor Pluggable (SFP) modules. We therefore require a separate dedicated input for clock distribution. The PAT cards are designed to be able to interface with the DUNE timing system [11] over an optical fibre. In order to implement this, extra optical feedthroughs are required to receive all the timing fibres into the pressure vessel. The number of feedthroughs can be mitigated by fanning out the signal inside the vessel but there are limits to the number of boards that can be supplied from any one fibre. In both cases, the clock received at the PAT cards in the pressure vessel is transmitted from the FPGA to a dedicated clock synthesizer for cleaning and then further transmitted to the FECs.
Furthermore, distribution of the FEC clock, sync and slow control signals happens, depending on the use case, via fan-out chips or multiplexers. Given that the designed readout is synchronous to the DUNE clock, all the FEC cards are expected to receive the same copy of the clock sync signal. For this reason, an organisation of eight FECs in seven groups have been designed. The advantage of this method relies on the layout design of the PCB. The chosen FPGA can drive seven individual groups with the same clock sync signal pulse. Each of these seven individual groups is then copied in phase to all eight FECs connected to that pin. The main reason for deciding to fan-out 1:8 was to ensure that the clock quality provided to all the FECs wasn’t impacted by the fan-out ASIC. Equivalently, for the slow control interface, a series of multiplexers is implemented allowing individual access to all the FECs, providing enhanced telemetry of the FEC operation.
Finally, the PAT cards are mounted at 3 positions on the OROC, in such a way that when all 18 OROCs are mounted into a circle, the PAT cards are shared between neighbouring OROCs. In this way, the cable lengths required between FECs and PAT cards are minimised and so, these connections are optimised. The PAT cards are mounted on aluminium plates, which, through thermal contact with the Read-Out Chamber (ROC), provide some amount of passive cooling.
2.2 Firmware design
The PAT card’s main processing unit is a Kintex Ultrascale FPGA package. This device was selected as it provides the required number of serial transceivers to accommodate both readout schemes for ND-GAr whilst satisfying the remaining requirements. The power consumption (based on the AMD power estimate tool) is within the nominal limits for operation in gases at atmospheric pressure or above when actively cooled. Finally, the availability of this package at low cost through CERN and the extensive experience of using it that has been built up at CERN make this FPGA a favourable choice given the cost driven nature of the design.
In order to take advantage of the FPGA, dedicated firmware had to be developed that allowed data to be reliably and efficiently streamed. Here, we discuss the firmware solution for direct powering, as this was the firmware that was eventually tested with the TOAD test beam. The design has two primary functions. First, the data flow, i.e. how FEC data are reliably received into the software. The second is Control Configuration and Monitoring (CCM) operations. A schematic representation of the firmware design can be seen in figure 4.
2.2.1 Data flow
The main objective of the data flow firmware is to reliably stream the FEC data into a server running the DUNE DAQ software [12]. The input to this firmware is the data stream from the daisy-chained SAMPAs, which is fed into the FPGA pins via a Low-Voltage Differential Signalling (LVDS) stream. The SAMPAs stream first a header which includes important information, like the number of samples that will be streamed, the time of arrival information, the fired channel number, etc. After the header, the payload follows which is made up of N words, where N is defined in the header.
The data stream is then decoded using dedicated decoder firmware based on a Finite State Machine (FSM) that can handle all the different operational modes of the SAMPA. The extracted payload along with some of the header information (channel number, timestamp, and number of clusters), is converted into a stream that is subsequently sent through a RR based buffering system. The output Ethernet sample contains additionally a single header with some of the header information received from the SAMPA header, as well as the DUNE timestamp for data alignment.
The firmware implements three options for the streaming protocol used to do this. The first option, which is the one we used for all data shown here, uses the standard IPbus library over User Datagram Protocol (UDP) [13]. This option provides a stream of a given number of words on request from the DAQ software.
The second option interfaces the SiTCP suite from BeeBeans [14] with IPbus. SiTCP is a lightweight interface that provides streaming capabilities over Transmission Control Protocol (TCP). We developed an interface between the standard pull-request based IPbus firmware and the SiTCP IP core, thereby providing a TCP based interface for the IPbus suite. While SiTCP includes its own UDP based interface for CCM, it does not have enough address space for our application.
The third option was used for the power over serial powering mode investigations. The number of serial links on the PAT only allows for a single hardware interface between the in-pressure and out-of-pressure boards for both data and CCM and hence the SiTCP solution can be used in the out-of-pressure electronics. The link between the in-pressure and out-of-pressure electronics still cannot use the SiTCP solution, so it is implemented as an Aurora based stream interface. In this operation mode, IPbus is not used at all and the received data are converted into a TCP stream interface using SiTCP.
2.2.2 Control, Configuration and Monitoring
The firmware includes many registers containing information allowing full telemetry of the hardware from software. The CCM aspect of the firmware’s job is to monitor and set these registers. The design of the PAT cards allows the FPGA to also monitor and change the behaviour of several other components on the board using standardised slow control interfaces (e.g. Inter-Integrated Circuit (I2C), Serial Peripheral Interface (SPI), Power-Management Bus (PMbus)). The list of components can be found in table 5 along with a short explanation of their function.
| Component | Interface | Application |
| FEC DC/DC Converter | PMbus | Power converter for FEC |
| SAMPA slow control | I2C | Configuration of SAMPA device |
| SFP Connectors | I2C | Ethernet I/O |
| Clock Synthesizer | I2C | Clock cleaning |
| Flash Memory | SPI | Remote firmware programming |
| EEPROM | I2C | Design information |
| I2C Switches | I2C | Slow control routing |
| Port-Expanders | I2C | Power switches for FEC |
Each detector operation mode requires a specific configuration of these components. For example, with the inclusion of the I2C port-expanders, individual FECs can be turned off and removed from the run if there is a problem with their operation.
In addition to the external hardware components, the CCM part of the firmware implements a variety of standard IPbus registers and memories dedicated to monitoring the data flow or configuring operational modes. A list of these registers can be seen in table 6.
| Reg. Name | Control/Monitor | Number of Registers |
| Clock control | Enable clock transmission to fanout chips | 1/FEC group |
| Clock rate control | Control over internal clock multiplexers for (, or ) | 1/PAT |
| Serial links control | Ethernet serial link monitor and configuration | 1/Serial Link |
| Sync. control | Send synch signal to all FEC sampling windows | 1/FEC group |
| Memory content | Live data flow memory content | 1/FEC |
| RR control | Enable, Veto and Reset of RR logic | 1/RR block |
| Logic Reset | Reset signals for all logic components | 1/Firmware block in the design |
2.3 Software development
Operating the PAT-based readout requires software as well as the custom firmware. The PAT cards have three main functions: data flow, CCM, and timing. Each function has a software counterpart that has been developed for the TOAD test beam and designed so as to minimise modifications needed for integration in the final ND-GAr. All three sets of software are designed to run on standard commercial servers.
The data flow application is based on the DUNE DAQ software suite. For CCM, an IPbus based framework has been developed allowing full access to all the in-pressure electronic’s configuration registers. Finally, for timing, the software is based on the framework from the DUNE timing system which we have wrapped in the bespoke PAT CCM framework since the timing system uses IPbus for configuration as well.
2.3.1 DUNE DAQ integration
Once the data has been streamed out, it is received by the DUNE DAQ software. While this software already existed prior to this work, significant effort was required to integrate TOAD’s DAQ into the larger DUNE software stack. A schematic of the general data flow within DUNE DAQ is shown in figure 6. The integration of the TOAD readout into this framework involved adding the Ethernet packet structure described in section 2.2.1 into the upstream DAQ. Additionally, the zero-suppression functionality of the FECs allows for the PAT boards to only output data above the threshold we wish to be considered for offline processing. This means that TOAD also demonstrated that the DUNE DAQ software framework worked in untriggered continuous readout mode.
A further difference between TOAD and the other systems using DUNE DAQ is that the output stream data format allows each hit that is streamed out to contain a variable number of ADC samples, while all other systems had fixed numbers of samples per transmitted data packet. This implementation expands the capabilities of the DUNE DAQ software and allows it to handle variable length data packets[12]. We carried out a test of this implementation and verified that it worked using real PAT boards and data from simulated muon tracks in an HPgTPC.
2.3.2 Control, Configuration and Monitoring software suite
The data from TOAD is received by standard DUNE software [12]. However, the CCM software suite is a new scalable and modular software framework. The framework has two key base classes to talk to the two classes of objects on the PAT boards: firmware-implemented registers and external devices. The framework is summarised in figure 7. Each base class contains general methods needed by all registers of that type. Communication to the firmware registers is via IPBus, so the base class contains IPBus specific methods, while the inherited classes for each specific register contain their own specific bit maps that relate their 32-bits to specific functions. This scheme allows scalability to all electronic devices that use IPBus and similar firmware implementation.
The external device base class contains methods that generate instructions for the FPGA to access registers on external devices. For a device that uses the I2C interface, the inherited class will contain an additional I2C address and a list of register addresses for the different registers inside the device. Additional slow control protocols, such as SPI or PMbus, are integrated in a similar manner.
The CCM suite also provides various low-level scripts that act as a Command Line Interface (CLI) for the user to manage the hardware. By combining multiple low-level scripts, one can create high-level scripts that are tailored towards specific use cases. Examples include the noise measurements outlined in section 4.1 and continuous temperature and power monitoring used in section 4.2.
3 Test environment
3.1 TOAD
The TOAD detector design and operation largely follow that of the UK high-pressure platform in [9]. Our amplification wire chamber comes from the ALICE [16, 6] at the CERN LHC, which studies quark-gluon plasma in high-energy Pb-Pb collisions. The ALICE TPC is the main component of the experiment for tracking and particle identification, requiring the ability to track particle densities of up to per unit of rapidity at , with a maximum trigger rate of [8]. The TPC’s shape is that of a hollow cylinder with a length of and radius of , leaving space in the centre for ALICE’s cylindrical inner tracking system with a radius of . The readout plane at each endplate is segmented into 18 trapezoidal sectors, with an Inner Read-Out Chamber (IROC) and an OROC per sector. Each ROC is a MWPC. Given the increased Pb-Pb luminosity delivered to the LHC in Run 3, corresponding to a collision rate of , the ALICE TPC required an upgrade to avoid pile-up [7]. Hence, the ROCs can be repurposed and an OROC is used as the readout chamber for the TOAD TPC.
3.2 Pressure vessel
TOAD utilises the same pressure vessel described in [9, 17], which allows the testing of the readout electronics with an ALICE OROC at pressures up to (). The vessel has an inner diameter of and an outer diameter of , made from Stainless Steel. The vessel has domed ends, giving a total volume of . The vessel can be opened by removing one of the domed faces. The vessel door is secured to the vessel body via a DN 1500 flange, where the flange on the vessel body has two grooves for O-rings: a diameter circular viton inner ring, and a rectangular silicon outer ring. The inner ring ensures chamber pressure tightness in vacuum conditions and the outer ring ensures it in high-pressure conditions. A detailed schematic of the vessel and its dimensions is shown in figure 8 and an image of the vessel in the test environment is shown in figure 9. There are five DN200 flanges on the vessel door, which are not used. One DN200 flange on the back of the vessel is used to mount a camera via a thick quartz optical window. There are various KF25 and KF40 flanges on the vessel, which are used for the High Voltage (HV), gas, vacuum and readout-electronics systems feedthroughs. The gas and evacuation system is connected to the vessel so that the vessel can be evacuated to . This is achieved by first evacuating with a roughing pump, and once the vessel is below , evacuation with a turbo pump through a Carten solenoid valve. The pump system used is an Edwards Turbo Cart which consists of a backing pump, turbo pump and controller. The pressure inside the vessel is monitored using two pressure sensors - an Inficon PGC550 combined capacitance-Pirani vacuum gauge for below-atmosphere pressure readings and a Wika A-10 digital pressure gauge for above-atmosphere pressure readings ( to ). These are attached directly to the vessel via feedthroughs and readings are stored regularly through the slow control system. The gas system allows the vessel to be filled with up to of gas mixtures, with the gas mix controlled by filling to a given partial pressure of gas from up to four different gas cylinders connected to the system. For the run described here, we were able to mix the following three input gases: premixed Ar-CH4 with molar ratio 92:8 for Ar:CH4, pure Ar, and pure N2.
The TPC consists of a HV drift cathode, which is a steel mesh ring of radius , made with wires of diameter with 25 lines per inch. There are three HV feedthroughs on the vessel, which are used to apply voltages to the cathode, anode and gating grid. Other feedthroughs are also used to ensure the equipment inside the vessel is all safely grounded. The cathode and OROC anode have been tested with voltages up to 16 kV and 3 kV respectively, with the maximum achievable voltage varying with the gas mixture used [9]. The field cage consists of eight copper rings of inner diameter , connected by resistors, with resistors of this resistance also connecting the first ring to the drift cathode and the last ring to ground [17]. The results shown in this paper use various configurations of HV, gas mixtures and pressures. These are summarised in table 7.
| Mode | Gas Mixture | Pressure (barA) | Anode (V) | Gating Grid (V) | Cathode (V) |
| 1 | Air | 1.0 | 0 | 0 | 0 |
| 2 | Ar:CH4, 96:4 | 4.5 | 0 | 0 | 0 |
| 3 | Ar:CH4, 98:2 | 3.0 | 1500.0 | -100.0 | 0 |
| 4 | Ar:CH4, 98:2 | 3.0 | 1500.0 | -100.0 |
In addition to the connections described above, the readout electronics were powered and data streamed through KF40 feedthroughs on the vessel door. One feedthrough was used to provide power to all three PAT boards. Each board then communicates through an individual RJ45 feedthrough connected directly to a COTS network switch which is, in turn, connected to two DELL PowerEdge servers: one running CCM software, and the other DAQ software – both of which are outlined in detail in section 2.3. Other feedthroughs on the vessel body are used for safety burst discs – a burst disc backed by a pressure relief valve and an independent burst disc direct to atmosphere – to ensure the vessel pressure does not exceed its rated pressure.
3.2.1 ALICE Multi-Wire Proportional Chambers
The pre-upgrade ALICE ROCs employ a typical wire plane schematic: an anode-wire grid above the pad plane, a cathode-wire grid and a gating grid to separate the drift volume from the readout. The pad plane and the three wire planes are all spaced apart. All wires run in the azimuthal direction (approximately vertically in the orientation we have the chamber) and have different separations in the various planes. The length of the shorter and longer parallel sides of the OROC are and , respectively. The perpendicular length of the OROC is . The gating grid wires have a pitch (spacing) of and are made of copper beryllium with a diameter. The cathode wires have the same properties but with a pitch of and are staggered to the gating grid wires. Likewise, the anode wires are staggered compared to the cathode and gating grid wires, with a pitch of . These anode wires, in contrast to the other two planes, must withstand a HV and are made from gold-plated tungsten with a diameter. The edge wires of the anode (first two and last two) have a thicker diameter of . A gating grid wire voltage of or is set to alternate wires. In this way, drift electrons or ions cannot pass the gating grid region. The cathode wires are held at ground. However, as the OROC used in TOAD was part of the ALICE ‘test’ ROCs, rather than a production ROC, the ability to change the bias at the cathode wires is a feature. The wires can be seen on the OROC in figure 10(a). The OROC has previously been successfully tested up to anode voltages of in Ar-CO2 and Ar-CH4 mixtures with pressures above [9]. The pad plane consists of 9984 pads in total, of two sizes, which are summarised in table 8. Here, “inner” refers to the rows closer to the shorter parallel side, whereas “outer” refers to rows closer to the longer side of the OROC. The difference in pad dimensions is due to the trapezoidal shape of the ROC. Groups of 21 or 22 pads are read out via a socket on the back of the OROC and the sockets are grouped in threes. As shown in figure 2, a TOAD FEC has three connectors, meaning that it can plug into three sockets and hence reads out 64 pads. The FECs and PAT boards are shown connected to the OROC in figure 10(b).
| Pad Dimensions (mm2) | Pad Rows | Number of Pads | |
| Outer Pads | 6 15 | 32 | 4032 |
| Inner Pads | 6 10 | 64 | 5952 |
| Total | - | 96 | 9984 |
In the TPC, ionisation electrons are produced and then drift towards the ROC due to the presence of an electric field, from the HV cathode electrode and the gating grid. When the electrons reach the ROC, they undergo avalanches at the anode wires and therefore, charge is induced onto the pads. The spatial position of the ionisation electrons in the ROC plane is determined from the pad signals. The perpendicular position, in the drift direction, is calculated from the comparative time of the signal on the ROC and a time, , which would come from beamline instrumentation triggers. The studies outlined in this paper are noise studies and, therefore, particle tracking is not employed.
4 Results
4.1 Noise characterisation
Understanding the noise profile is crucial in verifying that we have a sufficient signal-to-noise ratio to identify the signals we are looking for. We characterise noise as the Root Mean Square (RMS) error of the waveform in the absence of any signal.
We compute noise by taking the RMS of a waveform, which corresponds to 500 samples at , for each channel in each FEC. Figure 11 shows the spatial distribution of the RMS noise of all FECs in the OROC, measured under pressure. Channels with RMS noise less than are coloured grey, while channels with noise above are shown with the highest colour in the scale. The blank regions are caused by FECs that did not respond to any I2C communications.
Figure 12 plots the distribution of the RMS noise. Noise from a total of 9600 (9656) channels was measured for the atmospheric pressure () case, which corresponds to 96% of the total number of OROC pads as described in table 8. We require the waveform baseline to be greater than counts in figure 12. This cut is placed to remove waveforms that have noise cut off at due to the low baseline, which makes the waveform appear to have low noise. This cut removes of the data, and results in 8621 (8645) channels for the atmosphere () case. Correction of the baseline can be applied to individual SAMPA channels, which would mitigate the issue of low baseline and therefore recover these channels. This calibration was not performed during this operation period of TOAD due to time constraints.
The noise distribution is bimodal with a single-bin peak at just above 0 and another peak at . The first peak is caused by FECs that are not properly connected to the OROC. This peak is spatially shown as the grey patches in figure 11. This is demonstrated by measuring the RMS noise from the FECs after they are removed from the OROC and left floating. Figure 13 shows a comparison between the RMS noise for the same subset of 12 channels on one FEC, before and after the card is unplugged from the OROC. When the FEC is floating, all of the measured channels have RMS noise of . This level of noise corresponds to the intrinsic noise level of the SAMPA (), as shown in table 1, which is equivalent to and . A gain of is used for the conversion [10]. For most of the FECs that are well-connected to the OROC, their noise distribution is characterised by the second peak.
The large tail that extends past in figure 12 are shown as distinct noisy regions in figure 11. This is due to a partial connection between the FEC and the OROC which results in a floating ground, as this was not observed in benchtop tests with direct connection to the FECs. The difficulty in connecting the FEC to the OROC, for both the case of improper connection, which left the FEC afloat, and partial connection, which resulted in a floating ground, is due to the legacy OROC connectors used in TOAD. It is observed that the same FEC can gain good performance when moved to a difference slot on the backplane. It is also worth noting that the ALICE ROCs are now end of life, so a future HPgTPC would be designed so as to achieve a more reliable connection to the front-end electronics.
From figure 12, no noticeable difference can be seen in the noise measurements taken in atmosphere (mode 1 of table 7) and 96:4 Ar-CH4 (mode 2). A Kolmogorov-Smirnov (K-S) p-value of is obtained from the unbinned data, which suggests the same underlying distribution.
Additionally, we see that powering of the HV devices does not have a noticeable effect on the noise. This can be seen in figure 14 which compares the noise when the HV devices are off in atmosphere (mode 1 of table 7) and when in 98:2 Ar-CH4 (mode 3). A K-S p-value of is obtained and again suggests the same underlying distribution. As the data for this plot were taken in a different condition and before some noise improvements were made compared to figure 12, no direct comparison could be made. Nevertheless, the bimodal distribution and the positions of the peaks are still the same between the two sets of data. We also carried out a test with the same OROC and gating grid voltages but with the drift cathode powered to (mode 4), and the main features in the noise distribution remained identical.
The contribution of noise from different frequencies can be studied through Discrete Fourier Transformation (DFT) of the waveforms. Figure 15 shows the DFTs from a single channel on a FEC. To avoid random noise, 1000 waveforms are collected and the average of the DFT of 1000 individual waveforms is plotted. Large peaks at and integer multiples of it can be noticed in figure 15. This corresponds to the switching frequency of the FEC DC/DC converter. Less prominently, the smaller peaks at multiples of corresponds to the switching frequency of the FPGA DC/DC converter.
To summarise, the DC/DC components on the PAT boards are major contributors to the noise measured by the FECs. In addition, FECs that were not conductively connected to the OROC would appear to have extremely low noise. High noise channels in the OROC are seen in small regions, usually corresponding to a single FEC, and could be due to improper grounding of the OROC connectors.
To assess whether the level of noise measured in TOAD would be sufficient for a neutrino experiment, we estimate the size of the signal on a FEC that could be generated by a minimally ionising particle. The expected signal-to-noise ratio can be calculated using stopping power , atmospheric argon gas density , and ion pair production energy , giving at atmosphere[18]. At , using a projected gas gain of from [9] and accounting for the change in gas density, one obtains . We take as an approximate track length across an OROC pad, i.e., a channel in a FEC. Simulations have been done to show that the electrons spread across multiple pads, with only 22% lie within of the centre. Furthermore, it takes 7 samples for the charge to be fully deposited. So using from [10], on average, a minimally ionising particle is expected to deposit per channel per sample. The 95 percentile of figure 12 gives a conservative estimate for noise of , which leads to an estimated signal-to-noise ratio of 21.
4.2 Thermal profile
To achieve a reasonable gain in a gas TPC, a stable pressure must be maintained. The thermal profile of the system thus needs to be well-understood as in a closed vessel the temperature directly affects the pressure. Cooling is currently achieved in TOAD through thermal contact between the aluminium mounting plate of the PAT board and the OROC. The FECs are cooled through contact with the gas in the vessel. The temperature and the power output from each PAT board are recorded via the FEC DC/DC converter on each board, and the average values in are plotted in figure 16. The idle power consumption, with no power delivered to the FECs, for the PAT boards all connected to the same power bus ranges between 0.89-.
The FECs constitute the majority of the power consumption. With all available FECs in full operation, the power consumption for each PAT board ranged between , , and for the PAT board with , , and FECs connected, respectively. Assuming the power consumption scales linearly with the number of FECs, an average power consumption of per FEC can be calculated. This corresponds to a maximum power consumption of per PAT board.
Relying solely on the thermal coupling between the PAT boards and the OROC, an average change of temperature of on the DC/DC converter was observed as a result of the operation of the FECs. While the final temperature reached is within the operation threshold of the electronics, a large temperature change and use of the gas as the cooling medium can lead to a long stabilisation time for the pressure in the vessel as well as uncertain gas drift properties. As shown in figure 17, by fitting an exponential to the pressure profile, it takes for the pressure to have a rate of change of pressure . It is also worth noting that the heavy pressure vessel is in contact with the gas, increasing the time to heat up. Furthermore, the OROC is also in thermal contact with the TOAD vessel and the gas, further increasing the pressure stabilisation time and also exposing it to the high equilibrium temperature. Such a long stabilisation time prolongs the time taken to go from the detector being off to stable operations, increasing downtime.
Each PAT board is contributing a maximum of only of power which is well within the range that can be mitigated by the addition of active cooling or a better heat sink design. This investigation of the thermal profile is a necessary component of future research.
5 Conclusion
Many readout systems developed for particle physics experiments target the high-occupancy environment of a hadron collider experiment. We have described a system designed instead for much lower occupancy detectors, such as those in neutrino physics experiments. The system leverages the lower buffering requirements to use far fewer FPGAs per channel and thereby significantly reduce the cost of the system.
The ND-GAr detector that this system was designed for requires operation in a magnetic field and at high pressure (up to ). The high-pressure requirement also necessitates a limited number of cables exiting the pressure vessel. We have described the designed system and demonstrated its operation at a range of pressures and gas mixtures up to . These tests were carried out through the TOAD experiment and involved the readout of channels of an MWPC from the ALICE TPC.
The system performed well, stable operation was seen throughout the test period which included multiple days of running at atmosphere and pressure as well as bench top testing. However, the passive cooling used in the test has been identified as an area that must be improved before final deployment. Nevertheless, the total heat load is well within the range that can be dealt with active cooling. The noise characteristics of the system were measured and found to be within the range that will allow us to read out the intended signals for almost all channels. We have identified poor connection due to the mounting of the FECs in the legacy connectors of the OROC as the primary item contributing to channels that are not readable. This is specific to the OROC used in TOAD, and is not expected to be a problem for future HPgTPC e.g. ND-GAr which will use different connectors. Notwithstanding these necessary minor modifications, we have shown that this system is able to fulfil the needs of a high-pressure gas TPC detector for a fraction of the price of systems used for LHC type detectors.
Acknowledgments
We would like to thank the STFC, Royal Society and UKRI, UK; the DOE, USA; and the Imperial College President’s Scholarship scheme for funding this work. Particularly, for Fermilab, this document was prepared using the resources of the Fermi National Accelerator Laboratory (Fermilab), a U.S. Department of Energy, Office of Science, Office of High Energy Physics HEP User Facility. Fermilab is managed by Fermi Forward Discovery Group, LLC, acting under Contract No. 89243024CSC000002.
We would also like to acknowledge the invaluable assistance of the mechanical and electrical workshops at Royal Holloway University of London and Imperial College London, as well as the staff at the Fermilab test beam facility.